# Degrees of interior polynomials and parking function enumerators

### v2

@@ 2. The degree of the interior polynomial of a digraph
In our cases, $\tilde{\mathcal{Q}}_{G}$ is a $(|V|-1)$-dimensional polytope. Thus if we show that
@@
First, let us make a simple observation about dijoins. Let $C$ be any cycle in $G$. We call $C$ an signed cycle if we have picked one of its cyclic orientations as “positive.” Then we can partition $C$ to $C^{-}\sqcup C^{+}$ where $C^{-}$ is the set of edges pointing in the negative cyclic direction, and $C^{+}$ is the set of edges pointing in the positive cyclic direction. Signed cycles are not to be confused with directed cycles; the latter means a cycle whose edges all point in the same cyclic direction.
Suppose that $C$ is a signed cycle with a choice of orientation such that $|C^{-}|>|C^{+}|$. If a dijoin $K$ contains $C^{-}$, then $K$ is not minimal. Indeed, in this case $|K-C^{-}\cup C^{+}|<|K|$ and $K-C^{-}\cup C^{+}$ is also a dijoin. This is so because if a directed cut does not intersect $C$, then it is covered by $K-C^{-}\cup C^{+}$ the same way as by $K$; if a directed cut intersects $C$ then it intersects both $C^{-}$ and $C^{+}$, hence it is also covered by $K-C^{-}\cup C^{+}$.
Let us call a dijoin $K$ a long arc dijoin if it contains $C^{-}$ for some signed cycle $C$ with $|C^{-}|>|C^{+}|$. Our previous observation implies that long arc dijoins are never minimal. Note also that $C^{-}$ and $C^{+}$ are not meant to be actual arcs of $C$.
We claim the following characterization of relative interior points of ${\tilde{\mathcal{Q}}}_{G}$. The distinction between $<1$ and $\leq 1$ is a crucial subtlety in the statement. Note that since $\mathbf{0}$ is a generator of ${\tilde{\mathcal{Q}}}_{G}$, our formulas can be thought of as convex linear combinations even when the sum of the coefficients is less than $1$.
@@
A point $\mathbf{p}\in\tilde{\mathcal{Q}}_{G}$ is in the relative interior of ${\tilde{\mathcal{Q}}}_{G}$ if and only if there exists a cycle-free dijoin $K$ of $G$ such that $\mathbf{p}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$, where $\lambda_{e}>0$ for each $e\in K$ and $\sum_{e\in K}\lambda_{e}<1$.
@@
The following Lemma is equivalent to the well-known fact that ${\tilde{\mathcal{Q}}}_{F}$ is a unimodular simplex for any forest $F$. See, e.g., [13, Corollary 3.6] for a proof.
@@
For a forest $F$ and any positive integer $s$, a point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$ is a lattice point if and only if $\mathbf{p}=\sum_{e\in F}\mu_{e}\mathbf{x}_{e}$, where each $\mu_{e}$ are integer.
@@
Let $K$ be a dijoin of $G$ with cardinality $\nu(G)$. Then $\mathbf{p}=\sum_{e\in K}\mathbf{x}_{e}+\mathbf{0}$ is a point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$, moreover, it clearly has integer coordinates. Now by Proposition 2.10 we have that $\mathbf{q}=\frac{1}{\nu(G)+1}\mathbf{p}=\sum_{e\in K}\frac{1}{\nu(G)+1}\mathbf{x}_{e}$ is an interior point of $\tilde{\mathcal{Q}}_{G}$, which implies that $\mathbf{p}$ is also an interior point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$.
@@
The interior lattice points of $(\nu(G)+1)\cdot{\tilde{\mathcal{Q}}}_{G}$ are exactly the net degree vectors of the minimal dijoins of $G$.
@@
By Lemma 2.11, such a $\mathbf{q}$ is a lattice point if and only if each $\lambda_{e}$ is integer. Hence we conclude that $\lambda_{e}=1$ for each $e\in K$ and $|K|\leq\nu(G)$. Consequently $\mathbf{q}$ is an interior lattice point if and only if $K$ is a minimum cardinality dijoin, and $\mathbf{q}$ is the net degree vector of $K$. ∎
@@ 3. Degree-minimizing orientations
In [12] we looked at a special case of this problem, and considered the semi-balanced orientations of a bipartite graph $\mathcal{G}$ with partite classes $U$ and $W$. Any bipartite graph $\mathcal{G}$ has some (typically, many) semi-balanced orientations, but there are two special ones among them: The one where each edge is oriented from $U$ to $W$, and the one where each edge is oriented from $W$ to $U$. It is easy to see that these orientations are indeed semi-balanced. We call them the standard orientations of $\mathcal{G}$. The root polytopes of the two standard orientations are isometric, as they are reflections of each other. In particular, their interior polynomials coincide.
After looking at several examples, in [12] we conjectured that among all semi-balanced orientations of $\mathcal{G}$, the standard orientations minimize every coefficient of the interior polynomial. See [13, Example 6.5] for some concrete instances of this phenomenon. With Theorem 1.1 in hand, we are able to prove a weakened version of the conjecture, showing that the degree of the interior polynomial is minimized by the standard orientations. In fact, we can prove this among all orientations.
@@
Let $G$ be an arbitrary semi-balanced orientation of $\mathcal{G}$. By the Lucchesi–Younger theorem, the maximal number of disjoint directed cuts in $G$ is equal to $\nu(G)$, the minimal cardinality of a dijoin in $G$.
@@
On the other hand, [8, Theorem 9.6.12] claims that the standard orientations have $c(\mathcal{G})$ disjoint directed cuts. Hence they do maximize $\nu$ among all orientations of $\mathcal{G}$, as claimed. ∎
Concrete examples (e.g., [13, Example 6.5]) show that typically, there are some non-standard orientations that also minimize the degree of the interior polynomial. It would be interesting to give a characterization for all the other orientations that attain the minimal degree.
For non-bipartite graphs, we have no candidates for the degree-minimizing orientations. Yet, based on computer calculations, we make the following conjecture.
Conjecture 3.2.
For any undirected graph, there exists an orientation that coefficientwise minimizes the $h^{*}$-polynomial of the extended root polytope among all orientations. Also, there exists an orientation that coefficientwise maximizes the $h^{*}$-polynomial of the extended root polytope among all orientations.
@@ 4. A generalization to regular matroids
It turns out that many properties of the root polytopes of digraphs extend word-by-word to regular oriented matroids, moreover, in some applications, one needs this more general case (see, for example, Section 5 or [20]). Here we briefly recall the notion of a regular matroid, along with ways of orienting such a structure, and then generalize the results of Section 2 to this context.
From among the many equivalent characterizations of the class of regular matroids, we will use the following: A matroid is regular if it can be represented by the column vectors of a totally unimodular matrix. Here a matrix is totally unimodular if each of its subdeterminants is either 00, $-1$, or $1$.
More concretely, let $A$ be an $n\times m$ totally unimodular matrix with columns $\{\mathbf{a}_{i}\}_{i\in E}$. The ground set of our matroid is going to be $E$. A set of elements $C=\{i_{1},\dots,i_{s}\}\subseteq E$ is a circuit if the corresponding columns $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$ are minimally linearly dependent (over $\mathbb{R}$). Total unimodularity implies that in this case, the coefficients of the linear relation can be chosen from $\{-1,1\}$. For a circuit $C$, we call the two vectors in $\ker(A)\cap\{0,1,-1\}^{n}$, whose support is $C$, the vectors of $C$.
The orientation of the matroid is defined as follows. For each circuit $C$, any vector $\chi_{C}$ of $C$ gives a partition $C=C^{+}\sqcup C^{-}$, where $C^{+}$ is the subset of $C$ with positive coordinates in $\chi_{C}$, and $C^{-}$ is the subset of $C$ with negative coordinates. For different vectors of $C$, the subsets $C^{+}$ and $C^{-}$ might switch roles but the partition itself is well-defined because of the minimality of $C$. By a regular oriented matroid we mean a pair $(E,\mathcal{C})$, where $\mathcal{C}$ is the system of (unordered) pairs $(C^{+},C^{-})$ defined as above from a totally unimodular matrix $A$.
By a signed circuit, we mean a circuit $C$ with a choice of orientation, that is, with a choice of a vector of $C$ that determines which subset of $C$ should be $C^{+}$ and which one should be $C^{-}$.
Let us stress that a given regular oriented matroid may have several different totally unimodular representing matrices. (For example, row elimination steps performed on $A$ do not change the oriented matroid.)
One important example of regular oriented matroids is that of directed graphs. To a directed graph, one can associate a regular oriented matroid using its (directed) vertex-edge incidence matrix. The circuits of this matroid correspond to the cycles of the graph, and the partition $C=C^{+}\sqcup C^{-}$ of a circuit is to edges of the cycle pointing in the two cyclic directions.
@@
We call a regular oriented matroid co-Eulerian if $|C^{+}|=|C^{-}|$ for each circuit $C$. For graphic oriented matroids, being co-Eulerian is equivalent to the orientation of the graph being semi-balanced.
@@
It turns out that if $A$ and $A^{\prime}$ are two totally unimodular matrices representing the same oriented matroid $M$, then the $h^{*}$-polynomials of ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$ are the same [20]. (We note that [20] proves this for $\mathcal{Q}_{A}$ and $\mathcal{Q}_{A^{\prime}}$, but a straightforward modification of the argument yields the result also for ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$.) Hence the $h^{*}$-polynomial of ${\tilde{\mathcal{Q}}}_{A}$ is an invariant of the regular oriented matroid $M$, which we call the interior polynomial and denote by $I_{M}$. Note that the orientation of the matroid matters: if we keep the (unoriented) matroid structure, but change the orientation, then the interior polynomial might change. (This is true even for graphs, cf. [13, Example 6.5].)
@@
Let $M$ be a regular oriented matroid of rank $r$. Then the degree of $I_{M}$ is equal to $r-\nu(M)$, where
@@
Let $A\in\mathbb{R}^{n\times|E|}$ be a totally unimodular matrix, representing a co-Eulerian regular oriented matroid $M$. Let $R$ be the linear span of the columns $(\mathbf{a}_{i})_{i\in E}$ of $A$. Then the extended root polytope satisfies
@@
Next, we elucidate the connection of dijoins to the geometry of the extended root polytope. Analogously to the case of graphs, let us call a dijoin $K$ of a regular oriented matroid $M$ a long-arc dijoin if there is a signed circuit $C$ such that $|C^{-}|>|C^{+}|$ and $C^{-}\subseteq K$.
@@
In the proof, we will use the Farkas lemma of totally unimodular matrices (see for example [8, Lemma 4.2.13]). Let us state the lemma in the form that we will need. For two matrices $P\in\mathbb{R}^{n\times a}$ and $Q\in\mathbb{R}^{n\times b}$, we denote by $(P|Q)\in\mathbb{R}^{n\times(a+b)}$ the matrix obtained by writing the columns of $Q$ after the columns of $P$.
@@
Let $F$ be an independent set in the regular oriented matroid $M$, represented by the totally unimodular matrix $A$, and let $s\in\mathbb{Z}_{>0}$. A point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$ is a lattice point if and only if $\mathbf{p}=\sum_{i\in F}\mu_{i}\mathbf{a}_{i}$, where each $\mu_{i}$ is integer.
@@
Let $K$ be a dijoin of cardinality $\nu(M)$. Then $\mathbf{p}=\sum_{k\in K}\mathbf{a}_{k}+\mathbf{0}$ is a point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$, moreover, it has integer coordinates. By Proposition 4.2, $\mathbf{q}=\frac{1}{\nu(M)+1}\mathbf{p}=\sum_{k\in K}\frac{1}{\nu(M)+1}\mathbf{a}_{k}$ is an interior point of ${\tilde{\mathcal{Q}}}_{A}$. Hence $\mathbf{p}$ is also an interior point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$.
@@
The statement follows from Theorems 1.6 and 2.12, and Proposition 4.5. ∎
One may wonder if an analogue of Theorem 3.1 holds.
Problem 4.6.
Given a regular matroid, which orientation has the interior polynomial of smallest degree? Is there an orientation whose interior polynomial is coefficientwise minimal?

### v3

@@ 2. The degree of the interior polynomial of a digraph
Proof of Corollary 2.11.
This is a consequence of the following additions to the last two paragraphs of the previous proof. For any admissible layering $\ell$, the corresponding connected subgraph of $G$ contains a spanning tree and thus the supporting hyperplane $\{\mathbf{x}\mid\ell\cdot\mathbf{x}=1\}$ contains $|V|-1$ affine independent vertices of ${\tilde{\mathcal{Q}}}_{G}$. Hence we have a surjection from admissible layerings to facets not containing $\mathbf{0}$. Since the facet corresponding to $\ell$ has dimension $|V|-2$, its orthogonal complement with respect to $\mathbb{R}^{V}$ is $2$-dimensional and clearly spanned by $\ell$ and $\mathbf{1}$.
For an elementary directed cut $C^{*}$, the supporting hyperplane $\{\mathbf{x}\mid f_{C^{*}}(\mathbf{x})=0\}$ contains a facet through $\mathbf{0}$ because $E-C^{*}$ contains a two-component forest. Therefore the map given in (1) of the Corollary makes sense; it is surjective by the previous proof; finally it is also injective because $C^{*}$ can be recovered as the set of edges $e$ so that $\mathbf{x}_{e}$ is not in the facet. ∎
@@
In our cases, $\tilde{\mathcal{Q}}_{G}$ is a $(|V|-1)$-dimensional polytope by Proposition 2.6. Thus if we show that
@@
First, let us make a simple observation about dijoins. Let $C$ be any cycle in $G$. We call $C$ a signed cycle if we have picked one of its cyclic orientations as “positive.” Then we can partition $C$ to $C^{-}\sqcup C^{+}$ where $C^{-}$ is the set of edges pointing in the negative cyclic direction, and $C^{+}$ is the set of edges pointing in the positive cyclic direction. Signed cycles are not to be confused with directed cycles; the latter means a cycle whose edges all point in the same cyclic direction. We will need two simple lemmas on the relationship of cycles and minimal dijoins.
Lemma 2.13.
If $C$ is a cycle and $K$ is a dijoin such that $C\subseteq K$, then $K$ is not minimal (with respect to inclusion).
Proof.
Choose an orientation $C=C^{+}\sqcup C^{-}$ of $C$ so that $C^{-}\neq\emptyset$. We show that $K-C^{-}$ is also a dijoin. Indeed if a directed cut does not intersect $C$, then it is covered by $K-C$ and hence by $K-C^{-}$. Else if a directed cut intersects $C$, then it intersects both $C^{-}$ and $C^{+}$. Thus it is also covered by $K-C^{-}$. ∎
Lemma 2.14.
Suppose that $C$ is a signed cycle with a choice of orientation such that $|C^{-}|>|C^{+}|$. If a dijoin $K$ contains $C^{-}$, then $K$ is not a minimum cardinality dijoin.
Proof.
This follows from $|K-C^{-}\cup C^{+}|<|K|$ and $K-C^{-}\cup C^{+}$ also being a dijoin. That is so because if a directed cut does not intersect $C$, then it is covered by $K-C^{-}\cup C^{+}$ the same way as by $K$; if a directed cut intersects $C$ then it intersects both $C^{-}$ and $C^{+}$; in particular it is covered by $K-C^{-}\cup C^{+}$. ∎
Let us call a dijoin $K$ a long arc dijoin if it contains $C^{-}$ for some signed cycle $C$ with $|C^{-}|>|C^{+}|$. (Note that $C^{-}$ and $C^{+}$ need not be sequences of consecutive edges.) Lemma 2.14 implies that long arc dijoins never have minimum cardinality among dijoins.
We claim the following characterization of relative interior points of ${\tilde{\mathcal{Q}}}_{G}$. The distinction between $<1$ and $\leq 1$ is a crucial subtlety in the statement. Note that since $\mathbf{0}$ is a generator (although not always a vertex) of ${\tilde{\mathcal{Q}}}_{G}$, our formulas can be thought of as convex linear combinations even when the sum of the coefficients is less than $1$.
@@
A point $\mathbf{p}\in\tilde{\mathcal{Q}}_{G}$ is in the relative interior of ${\tilde{\mathcal{Q}}}_{G}$ if and only if there exists a dijoin $K$ of $G$, whose underlying undirected graph is cycle-free, such that $\mathbf{p}=\sum_{e\in K}\lambda_{e}\mathbf{x}_{e}$ with $\lambda_{e}>0$ for each $e\in K$ and $\sum_{e\in K}\lambda_{e}<1$.
@@
The following lemma is equivalent to the well-known fact that ${\tilde{\mathcal{Q}}}_{F}$ is a unimodular simplex for any forest $F$, cf. Lemma 2.2.
@@
Let us consider a forest $F$, a positive integer $s$, and a point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$. Then $\mathbf{p}$ is a lattice point if and only if $\mathbf{p}=\sum_{e\in F}\mu_{e}\mathbf{x}_{e}$, with each $\mu_{e}$ an integer.
@@
Let $K$ be a dijoin of $G$ with cardinality $\nu(G)$. By Lemma 2.13, this implies that the underlying undirected graph of $K$ is cycle-free. Then $\mathbf{p}=\sum_{e\in K}\mathbf{x}_{e}+\mathbf{0}$ is a point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$, moreover, it clearly has integer coordinates. Now by Proposition 2.15 we have that $\mathbf{q}=\frac{1}{\nu(G)+1}\mathbf{p}=\sum_{e\in K}\frac{1}{\nu(G)+1}\mathbf{x}_{e}$ is an interior point of $\tilde{\mathcal{Q}}_{G}$, which implies that $\mathbf{p}$ is also an interior point of $(\nu(G)+1)\cdot\tilde{\mathcal{Q}}_{G}$.
@@
The interior lattice points of $(\nu(G)+1)\cdot{\tilde{\mathcal{Q}}}_{G}$ are exactly the net degree vectors of the minimum cardinality dijoins of $G$.
@@
By Lemma 2.16, such a $\mathbf{q}$ is a lattice point if and only if each $\lambda_{e}$ is integer. Hence we conclude that $\lambda_{e}=1$ for each $e\in K$ and $|K|\leq\nu(G)$. By Lemma 2.13, minimum cardinality dijoins are all cycle-free. Consequently $\mathbf{q}$ is an interior lattice point if and only if $K$ is a minimum cardinality dijoin, and $\mathbf{q}$ is the net degree vector of $K$. ∎
@@ 3. Degree-minimizing orientations
In [14] we looked at a special case of this problem, and considered the semi-balanced orientations of a bipartite graph $\mathcal{G}$ with partite classes $U$ and $W$. Any bipartite graph $\mathcal{G}$ has some (typically, many) semi-balanced orientations, but there are two special ones among them: The one where each edge is oriented from $U$ to $W$, and the one where each edge is oriented from $W$ to $U$. It is easy to see that these orientations are indeed semi-balanced. We call them the standard orientations of $\mathcal{G}$. The root polytopes of the two standard orientations are isometric, as they are reflections of each other. In particular, their interior polynomials coincide. After looking at several examples, in [14] we suggested the following.
Conjecture 3.1.
Among all semi-balanced orientations of a bipartite graph, the standard orientations minimize every coefficient of the interior polynomial.
See [15, Example 6.5] for some concrete instances of this phenomenon. With Theorem 1.1 in hand, we are able to prove a weakened version of the conjecture, showing that the degree of the interior polynomial is minimized by the standard orientations. In fact, we can prove this among all orientations.
@@
Let $G$ be an arbitrary orientation of $\mathcal{G}$. By the Lucchesi–Younger theorem [18], the maximal number of disjoint directed cuts in $G$ is equal to $\nu(G)$, the minimal cardinality of a dijoin in $G$.
@@
On the other hand, [9, Theorem 9.6.12] claims that the standard orientations have $c(\mathcal{G})$ disjoint directed cuts. Hence they do maximize $\nu$ among all orientations of $\mathcal{G}$, as claimed. ∎
Concrete examples (e.g., [15, Example 6.5]) show that typically, there are some non-standard orientations that also minimize the degree of the interior polynomial. It would be interesting to give a characterization for all the other orientations that attain the minimal degree.
As to the largest possible degree of the interior polynomial, Theorem 1.1 readily implies the following.
Proposition 3.3.
Among all orientations of a connected graph $\mathcal{G}$, the degree of the interior polynomial is maximized by the following kinds of orientations: Consider the unique decomposition of $\mathcal{G}$ into a block sum of bridge edges (i.e., one-element cuts) and $2$-connected graphs. Orient the bridges arbitrarily and give each $2$-connected component a strongly connected orientation.
Proof.
Clearly, in any orientation, any bridge needs to be contained by any dijoin. We have a dijoin that contains only the bridge edges if and only if the $2$-connected components are strongly connected. ∎
Finally, we note that for general (non-bipartite) graphs, it can happen that there is no orientation that simultaneously minimizes all coefficients of the interior polynomial. Similarly, it can happen that there is no orientation that simultaneously maximizes all coefficients of the interior polynomial (the latter is possible also for bipartite graphs). We show two such instances in Example 3.4.
It would be interesting to understand which graphs have orientations that simultaneously maximize or simultaneously minimize all coefficients of the interior polynomial.
Example 3.4.
If the graph $\mathcal{G}$ is not bipartite, there is no notion of “standard” orientation, and there need not be any orientation coefficientwise minimizing the interior polynomial. For an example, consider the graph on the first two panels of Figure 2. Those two orientations have interior polynomials $12x^{3}+6x^{2}+4x+$ 1 and $x^{3}+7x^{2}+4x+1$, respectively, whence they are incomparable, and one can check that there is no orientation that would have an interior polynomial which is coefficientwise smaller than any of them.
It is also true that sometimes there is be no orientation coefficientwise maximizing the interior polynomial. See the graph on the $3^{\text{rd}}$ and $4^{\text{th}}$ panels of Figure 2. Up to isomorphism, there are $2$ strongly connected orientations here (one of them is on the $4^{\text{th}}$ panel), and they both have interior polynomial $x^{5}+3x^{4}+4x^{3}+4x^{2}+3x+1$. All other polynomials have 00 as the coefficient of $x^{5}$ by Theorem 1.1. Hence our only candidate for a coefficientwise maximal interior polynomial is the one above, but since the orientation on the $3^{\text{rd}}$ panel yields $12x^{4}+5x^{3}+5x^{2}+3x+$ 1, with larger coefficients for $x^{2}$ and $x^{3}$, the candidate fails.
[…]
Figure 2. Illustration for Example 3.4.
@@ 4. A generalization to regular matroids
It turns out that many properties of the root polytopes of digraphs extend word-by-word to oriented regular matroids, moreover, in some applications, one needs this more general case (see, for example, Section 5 or [22]). Here we briefly recall the notion of a regular matroid, along with ways of orienting such a structure, and then generalize the results of Section 2 to this context. For a more thorough introduction, see for example [5].
From among the many equivalent characterizations of the class of regular matroids, we will use the following: A matroid is regular if it can be represented by the column vectors of a totally unimodular matrix. Here a matrix is totally unimodular if each of its subdeterminants is either 00, $-1$, or $1$. More concretely, let $A$ be an $n\times m$ totally unimodular matrix with columns $\{\mathbf{a}_{i}\}_{i\in E}$. The ground set of our matroid is going to be $E$. The rank of a set $\{i_{1},\dots,i_{s}\}\subseteq E$ is defined as the rank of the submatrix formed by the columns $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$. A set of elements $C=\{i_{1},\dots,i_{s}\}\subseteq E$ is a circuit if the corresponding columns $\mathbf{a}_{i_{1}},\dots,\mathbf{a}_{i_{s}}$ are minimally linearly dependent (over $\mathbb{R}$). Total unimodularity implies that in this case, the coefficients of the linear relation can be chosen from $\{-1,1\}$. For a circuit $C$, we call the two vectors in $\ker(A)\cap\{0,1,-1\}^{n}$, whose support is $C$, the vectors of $C$. (Minimal linear dependence implies that there are only two such vectors.)
Each circuit $C$ supports two signed circuits, both being an ordered partition $C=C^{+}\sqcup C^{-}$, where for one of the vectors $\chi_{C}$ of the circuit, $C^{+}$ is the subset of $C$ with positive coordinates in $\chi_{C}$ and $C^{-}$ is the subset of $C$ with negative coordinates. For a given $C$, its two signed circuits differ by switching the roles of $C^{+}$ and $C^{-}$.
Let us stress that a given oriented regular matroid may have several different totally unimodular representing matrices. (For example, row elimination steps performed on $A$ do not change the oriented matroid.) On the other hand, if we multiply a column of $A$ by $-1$, the oriented matroid structure changes, while the unoriented matroid structure is preserved. (That is, the (unsigned) circuits remain the same, while the signed circuits change.) We note that for regular matroids, we do not lose any generality by concentrating on representing matrices; indeed, by [5, Corollary 7.9.4] all oriented matroid structures for a regular matroid can be obtained from a TU representing matrix.
One important example of oriented regular matroids is that of directed graphs. To a directed graph, one can associate an oriented regular matroid using its (directed) vertex-edge incidence matrix. The circuits of this matroid correspond to the cycles of the graph, and the partition $C=C^{+}\sqcup C^{-}$ of a circuit is to edges of the cycle pointing in the two cyclic directions.
@@
Signed circuits and signed cocircuits are orthogonal, which means that if a signed circuit $C$ and cocircuit $C^{*}$ have a nonempty intersection, then both $(C^{+}\cap(C^{*})^{+})\cup(C^{-}\cap(C^{*})-)$ and $(C^{+}\cap(C^{*})^{-})\cup(C^{-}\cap(C^{*})+)$ are nonempty [5, Theorem 3.4.3]. In particular, this means that if $C^{*}$ is a directed cocircuit, then if it intersects a circuit $C$, it intersects both $C^{+}$ and $C^{-}$.
We call an oriented regular matroid co-Eulerian if $|C^{+}|=|C^{-}|$ for each circuit $C$. For graphic oriented matroids, being co-Eulerian is equivalent to the orientation of the graph being semi-balanced.
@@
It turns out that if $A$ and $A^{\prime}$ are two totally unimodular matrices representing the same oriented matroid $M$, then the $h^{*}$-polynomials of ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$ are the same [22]. (We note that [22] proves this for $\mathcal{Q}_{A}$ and $\mathcal{Q}_{A^{\prime}}$, but a straightforward modification of the argument yields the result also for ${\tilde{\mathcal{Q}}}_{A}$ and ${\tilde{\mathcal{Q}}}_{A^{\prime}}$.) Hence the $h^{*}$-polynomial of ${\tilde{\mathcal{Q}}}_{A}$ is an invariant of the oriented regular matroid $M$, which we call the interior polynomial and denote by $I_{M}$. Note that the orientation of the matroid matters: if we keep the (unoriented) matroid structure, but change the orientation, then the interior polynomial might change. (This is true even for graphs, cf. [15, Example 6.5].)
@@
Let $M$ be a oriented regular matroid of rank $r$. Then the degree of $I_{M}$ is equal to $r-\nu(M)$, where
@@
Let $A\in\mathbb{R}^{n\times|E|}$ be a totally unimodular matrix, representing a co-Eulerian oriented regular matroid $M$. Let $R$ be the linear span of the columns $(\mathbf{a}_{i})_{i\in E}$ of $A$. Then the extended root polytope satisfies
@@
Next, we elucidate the connection of dijoins to the geometry of the extended root polytope. Analogously to the case of graphs, let us call a dijoin $K$ of a oriented regular matroid $M$ a long-arc dijoin if there is a signed circuit $C$ such that $|C^{-}|>|C^{+}|$ and $C^{-}\subseteq K$.
@@
In the proof, we will use the Farkas lemma of totally unimodular matrices (see for example [9, Lemma 4.2.13]). Let us state the lemma in the form that we will need. For two matrices $P\in\mathbb{R}^{n\times a}$ and $Q\in\mathbb{R}^{n\times b}$, we denote by $(P|Q)\in\mathbb{R}^{n\times(a+b)}$ the matrix obtained by writing the columns of $Q$ after the columns of $P$.
@@
We also need the following lemma, whose proof is a literal generalization of the graph case (Lemma 2.13), which is why we omit it.
@@
If $C$ is a circuit and $K$ is a dijoin such that $C\subseteq K$, then $K$ is not minimal (with respect to inclusion).
Lemma 4.5.
Let $F$ be an independent set in the oriented regular matroid $M$, represented by the totally unimodular matrix $A$, and let $s\in\mathbb{Z}_{>0}$. A point $\mathbf{p}\in s\cdot{\tilde{\mathcal{Q}}}_{F}$ is a lattice point if and only if $\mathbf{p}=\sum_{i\in F}\mu_{i}\mathbf{a}_{i}$, where each $\mu_{i}$ is integer.
@@
Let $K$ be a dijoin of cardinality $\nu(M)$. By Lemma 4.4, $K$ is independent in $M$. Then $\mathbf{p}=\sum_{k\in K}\mathbf{a}_{k}+\mathbf{0}$ is a point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$, moreover, it has integer coordinates. By Proposition 4.2, $\mathbf{q}=\frac{1}{\nu(M)+1}\mathbf{p}=\sum_{k\in K}\frac{1}{\nu(M)+1}\mathbf{a}_{k}$ is an interior point of ${\tilde{\mathcal{Q}}}_{A}$. Hence $\mathbf{p}$ is also an interior point of $(\nu(M)+1)\cdot{\tilde{\mathcal{Q}}}_{A}$.
@@
The statement follows from Theorems 1.7 and 2.17, and Proposition 4.6. ∎
One may wonder about the relationship of different orientations, in particular whether some analogue of Theorem 3.2 holds.
Definition 4.7.
A matroid is called bipartite if each circuit in it has even cardinality.
Problem 4.8.
Given a regular matroid, or a bipartite regular matroid, which orientation has the interior polynomial of smallest degree? Within the class of bipartite regular matroids, is there an orientation whose interior polynomial is coefficientwise minimal?